Ask your own question, for FREE!
Algebra 13 Online
OpenStudy (anonymous):

A game is played where 26 tiles are in a bag. Each tile has a different letter of the alphabet on it. If you draw four letters, what is the probability of the letters being M-A-T-H in order, if there is no replacement?

OpenStudy (anonymous):

1/26*1/25*1/24*1/23=1/358800=.000002787

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!